The USA and Australia are competing in the swimming medley relay during the Olympic Games. At the start of the final 50 meter stretch of the race, Eamon Sullivan pushes off the wall 2 seconds ahead of Michael Phelps. Sullivan swims at an average velocity of 2 meters/second, and Phelps swims at an average velocity of 2.18 meters/second. State which team won the race, how many seconds ahead of the second place swimmer the race winner finished, and how many meters ahead of the second place swimmer the race winner finished.
Answer by macston(5194)

.
Sullivan took (50m/2m/s)=25 seconds to finish.
Phelps took (50m/2.18m/s+2s)=24.936 seconds to finish.
Phelps won, finishing 0.064 seconds ahead of Sullivan.
Sullivan was (0.064s)(2m/s)=0.128 meters behind